4x+2y=4 What is the slope
2x-4y=10 What is the slope
Are the line paraller, perpendicular or neither


Answer by checkley77(12844)

Y=mX+b where the m value is the slope.
4x+2y=4
2y=-4x+4
y=-4x/2+4/2
y=-2x+2 thus this slope=-2
---------------------------------
2x-4y=10
-4y=-2x+10
y=-2x/-4+10/-4
y=x/2-2.5 thus this slope=1/2 or .5